# The Value of Information in Human-AI Decision-making

### v4

@@ Results.
Posterior predictions of human-AI team performance according to our regression model also align with our expectations based on the complementary information each model provides (Figure 1 far right). Collapsing across explanation conditions, AI1 reduces APE by more than AI2 ($6.27\%[5.91\%,6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively).
@@
Figure 1 also illustrates the effect of the different explanation conditions on the human-AI team performance. When the AI has sufficient complementing information (i.e., AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation, all for model AI1).
@@ 6 Discussion and Limitations
Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions. The basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021), even if human agents deviate considerably from Bayesian rational decision-making. We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making. Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy. Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023, 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
One direction for follow-up work is to develop tests that complement the framework by helping evaluators confirm the information that agents are using. Note that low ACIV of a signal over human decisions is not necessarily sufficient for concluding that particular signals were used by the human. It is often unknown whether a human has other private signals offering equivalent information to a feature being analyzed.
@@ Appendix A The Combinatorial Nature of the Value of Signals
10:  $output\ \phi_{j^{*}}$
Appendix B Estimating the Data-Generating Process from Observations
Suppose the data-generating process is estimated from the empirical distribution of the observed realizations of the payoff state value and the signal, $D=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
$\hat{\pi}_{D}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
In applying our framework, the rational DM’s decision rule $d^{r}$ should not overfit to the observed dataset. This threatens the usefulness of analysis using the DGP. To avoid this situation, we can either collect more observations or “coarsen” the signal space, where “coarsen” means reducing the signal space by aggregating similar signals into a single signal. We propose a systematic procedure for the latter since it may be too costly or otherwise not possible to collect more observations. Our objective is to find the coarsened signal structure that has the largest signal space and avoids the rational decision rule from overlapping. Denote the maximum tolerance factor for overfitting as $\epsilon$. The rational DM’s decision rule $d^{r}_{\hat{\pi}_{D^{\prime}}}$ that is estimated from a set of observations $D^{\prime}$ should perform similarly on $D^{\prime}$ and on the evaluation of a holdout set $D\setminus D^{\prime}$, i.e.,
$\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]-\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D\setminus D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]\leq\epsilon*\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]$ (12)
Algorithm 2 Estimating the data-generating process from a set of observations
1:  Input: Observed dataset $C=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Overfitting tolerance $\epsilon$, Ratio of the holdout set $\delta$.
2:  randomly split the dataset into a training set $C^{\prime}$ and a holdout set $C\setminus C^{\prime}$ with
$\lvert C^{\prime}\rvert=(1-\delta)*T$
3:  $\hat{\pi}_{C^{\prime}}(\omega,v)=\frac{1}{T}\sum_{\hat{\omega},\hat{v}\in C^{\prime}}\mathbb{1}_{\omega=\hat{\omega}}\cdot\mathbb{1}_{v=\hat{v}}$
4:  $R_{training}=\text{\bf E}_{(v,\omega)\sim C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
5:  $R_{testing}=\text{\bf E}_{(v,\omega)\sim C\setminus C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
6:  $bin\_size=\text{the number of unique }v$
7:  if $R_{training}-R_{testing}>\epsilon*R_{training}$ then
8:     $bin\_size=bin\_size/2$
9:     $\{v^{\prime}_{i}\}_{i=1}^{T}=\text{coarsen }\{v_{i}\}_{i=1}^{T}\text{ with binsize }bin\_size$.
10:     replace $\{v_{i}\}_{i=1}^{T}$ in $C$ with $\{v^{\prime}_{i}\}_{i=1}^{T}$.
11:     go to step 3.
12:  end if
13:  coarsened signal space $\mathbf{V}^{\prime}=\text{unique}(\{v^{\prime}_{i}\}_{i=1}^{T})$
14:  Output: $\hat{\pi}_{\pi^{C}}$, $\mathbf{V}^{\prime}$
Denote the optimal coarsened signal space that satisfies Equation 12 as
@@ Appendix B Estimating the Data-Generating Process from Observations
The above algorithm finds the coarsened signal space $\mathbf{V}^{\prime}$ that has the 2-factor approximation of the optimal coarsened signal space $\mathbf{V}^{*}$, $\lvert\mathbf{V}^{\prime}\rvert\geq\frac{1}{2}\lvert\mathbf{V}^{*}\rvert$, in $O(log(\lvert\mathbf{V}^{*}\rvert))$ time.
B.1 DGP with high-dimensional signals
The above algorithm’s time complexity is linear in the number of dimensions of the signal space. However, in practice, high-dimensional signals, such as images and text that appear in human-AI decision problems can still be computationally infeasible to quantify with the above algorithm. To address this issue, we propose a two-step approach to estimate the DGP from a set of observations. First, we use a reduction model to reduce the dimensionality of the signal space. Second, we use the estimated DGP using the reduced signal space in the above algorithm.
The optimal reduction model and its generated embeddings should fulfill two properties. First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals. Second, the derived DGP that is implied by the embedding signals should be feasible to estimate from the observed dataset. We quantify the above objectives by a similar algorithm based on Train/Test split.
@@ Results.
Appendix D Demonstration I: Model Comparison on Chest Radiograph Diagnosis
@@ D.1 Data and Model
We construct the scale of performances by a no-information bound and a full-information bound. The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value. The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models. We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals. We use the five vision models’ predictions as embeddings of the original images. See more detail on the algorithm of choosing the optimal embeddings in LABEL:app:high-dimensional.
D.2 Results
@@
Figure 3 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition 3.7, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
Appendix E Demonstration II: Behavioral Analysis on Deepfake Detection
@@ E.1 Data and Model
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. [2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Appendix D while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
E.2 Results
@@
Appendix F Demonstration III: Information-based Local Explanation on Recidivism Prediction
We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
Figure 5: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set. (C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18. The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output). See further comparisons between SHAP and ILIV-SHAP in Appendix H.
F.1 Data and Model
We use the COMPAS dataset [Angwin et al., 2022], which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. [2020] to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al. [2022]’s experiment on laypeople. Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model [Chen and Guestrin, 2016] on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. [2020].
We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism [Green and Chen, 2019a, Lin et al., 2020, Fogliato et al., 2021]. The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years. We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^6] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
F.2 Results
SHAP explanations do not communicate the complementary information of AI predictions.
Specifically, the changes in information value of AI predictions caused by marginalizing out features, i.e., ILIV-SHAP scores, are not predictable from the changes in AI predictions, i.e., SHAP scores, and the feature values. Figure 5 (A) and (B) compare the distribution of feature importance scores from SHAP to those from ILIV-SHAP. The scores of prior_counts (the number of prior criminal charges), age, and decile_score (the score from the COMPAS system), which are three features with the largest scores in both SHAP and ILIV-SHAP, show different correlations with the feature values in SHAP and ILIV-SHAP. For instance, in Figure 5(A), most of the low values (blue dots) of prior_counts have a negative SHAP score, and most of the high values (red dots) have a positive SHAP score. This indicates that when the value of prior_counts is low, marginalizing out prior_counts from the features leads to a higher expected model prediction than not marginalizing out, and vice versa. However, this correlation does not hold for ILIV-SHAP. In Figure 5(B), the high values of prior_counts appear on both the left side (below zero) and the right side (above zero), indicating having low ILIV-SHAP scores and high ILIV-SHAP scores at the same time. This mismatch between the distributions of SHAP scores and ILIV-SHAP scores implies that the impact of features on the information value cannot be directly inferred (at least monotonically) from their impact on model output and their values.
ILIV-SHAP can augment SHAP on communicating the complementary information of AI predictions.
We argue that ILIV-SHAP can supplement saliency-based explanations to provide users with a sense of how the AI predictions can help improve their decisions. With ILIV-SHAP, users can check whether the prediction contains decision-relevant information by looking at the total $\operatorname{ILIV}$ of the AI prediction, which indicates the maximum improvement in expected payoff that could be achieved by incorporating this prediction. Figure 5(D) and (F) show two different cases, where the AI predictions contain substantial human-complementary information value in (D) (shown by the distance between the tail and the head of the Waterfall plot) but very little information in (F).
ILIV-SHAP also tells the user whether an AI model’s prediction encodes useful complementary information from a particular feature. In Figure 5(E), the defendant’s prior count record (prior_counts=1) ranks second among features that impact the AI prediction ($-0.139$). Upon seeing this, users might be interested in investigating how the AI uses prior_counts=1. However, Figure 5(F) shows that prior_counts=1 contributes marginally to the decision-relevant information in AI predictions (about $+0.007$). Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.

### v5

@@ Results.
The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure 1 far right). Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.27\%$$5.91\%,$$6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively) [^5] See Appendix D for the significant test results.
@@
Figure 1 also illustrates the effect of the different explanation conditions on human-AI team performance. When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation).
6 Demonstrations
6.1 Chest X-ray Diagnosis
We study a chest X-ray diagnosis task. As agent decisions, $D^{\text{AI}}$ and $D^{\text{H}}$, we consider five predictive models fine-tuned on the MIMIC-CXR database (Johnson et al., 2019) and radiologists’ textual reports recorded in MIMIC-CXR. We use five pretrained image models with the same choices in Irvin et al. (2019). We train the models on 12,228 radiographs, and validated on 6,115 randomly sampled radiographs. For the payoff-relevant state, $\omega\in\mathbf{\Omega}=\{0,1\}$, we used results from two types of blood tests (‘NT-proBNP’ and ‘troponin’ cut by age-specific thresholds). The decision task is formalized as a prediction problem with $\mathbf{D}=\Delta\mathbf{\Omega}$ and $S(d,\omega)=1-(d-\omega)^{2}$.
Can the AI models complement human judgment?
Figure 2(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
Figure 2(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix F and Figure 6.
(a) Chest X-ray diagnosis results.
(b) Deepfake detection results
Figure 2: Demonstration results for both tasks. Distributions show information values. For a), we plot ACIV of radiologist decisions, different choices of AI models and signals that combine AI predictions and radiologist decisions. For b), we plot the information value of the combination of video-level features and agent decisions (including human decisions, AI predictions, and human-AI teams’ decisions). These two figures demonstrate that our framework can be used to a) reveal the complementary information between AI predictions and human decisions and b) analyze how humans use information in data features differently from AI.
As shown in this demonstration, doctors may use our framework to learn how much complementary information value the AI models offer over their decisions, and which model offers the most. [^6] For space constraints, we refer to the full descriptions and results in Appendices F and E.
6.2 Deepfake Detection
We study a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022). We select the model in the Deepfake Detection Challenge (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data. We use the human decisions and the human-AI team’s decisions from Groh et al. (2022), who collected judgments on the videos from n=5,524 participants recruited on Prolific. We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from Groh et al. (2022)’s use of mean absolute error, but again we prefer the quadratic score because it is a proper scoring rule where truthfully reporting beliefs maximizes the score. Groh et al. (2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm 1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin. We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure 2(b).
How much decision-relevant information do agents’ decisions offer?
We first compare the information value of the AI predictions to that of the human decisions. Figure 2(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$).
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different video-level features over different agents. As shown on the fifth row in Figure 2(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 2(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions.
@@ 7 Discussion and Limitations
Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions. Importantly, the basis of our framework in Bayesian decision theory does not require that actual (e.g., human) decision-makers achieve Bayesian rational decision-making. Rather, it provides a theoretical basis to support comparisons to human behavior to drive learning (see, e.g., (Guo et al., 2024; Hullman and Gelman, 2021; Wu et al., 2023). This theoretical basis is necessary to establish well-defined benchmarks. We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making. Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy. This suggests that valuing information value in terms of what it says about the payoff relevant state, not just the AI prediction, can improve the design of signals for human-AI decision-making. Hence our work offers theoretical support for attempts to design new explanations (e.g., Li et al. (2025)).
Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023, 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments. Though most of our definitions and analysis are focused on decision tasks with single well-defined payoff function, the framework can be readily extended to more complex decision tasks with an ambiguous set of payoff functions or totally unidentifiable payoff functions. We present a robustness analysis framework in Appendix B to calculate the Blackwell ordering of models over proper scoring rules. A demonstration of the robustness analysis is shown in Appendix F.
@@ Appendix A The Combinatorial Nature of the Value of Signals
Appendix B Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients [Mclaughlin and Spiess, 2023]. Blackwell’s comparison of signals [Blackwell et al., 1951] is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [Li et al., 2022, Kleinberg et al., 2023].
Definition B.1 (Blackwell Order of Information).
$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
where $\mathrm{R}^{\pi,S}(V)$ denotes the expected performance of the rational DM on payoff function $S$ and information structure $\pi$ when observing $V$.
Definition B.2.
Proposition B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
Proposition B.3 (Hu and Wu 2024).
If $\forall\mu\in(0,1)$
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
then $V_{1}$ is Blackwell more informative than $V_{2}$.
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix E.
@@ Results.
Appendix D Statistical Tests on Experiment Results
We calculate the p-values of the significance tests with Welch’s t-test, with $\alpha=0.05$. The degrees of freedom ($\nu$) of Welch’s t-test is calculated with the following formula:
$\nu=\frac{(\frac{s_{1}^{2}}{N_{1}}+\frac{s_{2}^{2}}{N_{2}})^{2}}{\frac{s_{1}^{2}}{N_{1}^{2}(N_{1}-1)}+\frac{s_{2}^{2}}{N_{2}^{2}(N_{2}-1)}}$ (10)
where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table 3). We used Bonferroni correction as the multiple comparisons correction to control the overall type I error rate. The results are shown in Table 2.
Table 2: 95% confidence intervals and significance test results (p-values) for the reduction of APE with different AI models and explanations in Figure 1.
P(ILIV-SHAP + SHAP
¿ SHAP)
P(ILIV-SHAP + SHAP
¿ No explanation)
P(AI1 ¿ AI2)
P(AI1 + ILIV-SHAP + SHAP
¿ AI2 + ILIV-SHAP + SHAP)
Reduction of APE with 95% confidence intervals
No explanation SHAP ILIV-SHAP + SHAP
AI1
$5.95\%$
$5.49\%,6.40\%]$
$6.06\%$
$5.65\%,6.47\%]$
$7.14\%$
$6.70\%,7.58\%]$
0.003 / 0.001
0.311 / ¡0.001
AI2
$6.24\%$
$5.79\%,6.68\%]$
$6.29\%$
$5.80\%,6.78\%]$
$5.47\%$
$4.95\%,5.98\%]$
0.138 / 0.159
Table 3: Sample sizes and standard deviations of the reduction of APE for each experimental condition
Model Explanation Participants Observations (N) Standard Deviation
AI1 ILIV-SHAP + SHAP 70 840 0.0653
SHAP 79 948 0.0642
No explanation 69 828 0.0668
Total (all explanations) 218 2,616 0.0656
AI2 ILIV-SHAP + SHAP 67 804 0.0746
SHAP 62 744 0.0679
No explanation 74 888 0.0673
Total (all explanations) 203 2,436 0.0700
Total 421 5,052 0.0678
Appendix E Demonstration I: Model Comparison on Chest Radiograph Diagnosis
@@ E.1 Data and Model
We construct the scale of performances by a no-information bound and a full-information bound. The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value. The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models.
E.2 Results
@@
Figure 4 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
Appendix F Demonstration II: Behavioral Analysis on Deepfake Detection
@@ F.1 Data and Model
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. [2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Appendix E while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
F.2 Results